OpenStudy (anonymous):

Input the equation of the given line in standard form. The line including (3, 1) and (-2, 3).

OpenStudy (anonymous):

I doing the same exact thing ughhh idk how to do it.

OpenStudy (anonymous):

I'm*

OpenStudy (anonymous):

me neither

OpenStudy (anonymous):

The line including A(3, 1) and B(–2, 3) slope AB is (3 – 1) ÷ (–2 – 3) = –2/5 = m Equation of line is Y – Ya = m(X – Xa) ➞ y –1 = –2/5(x – 1) ➞ 5y – 5 = –2x + 2 ➞ 2x + 5y = 7 in standard form (Ax + By = C)

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

actually it Xa should be 3

OpenStudy (anonymous):

so it should come out to 2x+5y=11